find the equation of the straight line which cuts off an intercept -5 from the y-axis and makes an angle of sin^-1 12/13 with the x-axis

Grade:11

1 Answers

Venky
7 years ago
given sin@ =12/13   , so tan @=12/5 the equation  y-Y =m(x-X)  given =(X,Y)=(0,5)
so y-5=12/5(x-0)     = 12x-5y+25